OpenStudy (anonymous):

Please help explain this to me. Rational Equations The equation is 100/45 + 75/35 = 150/t LCD: 20/9 + 15/7 =150/t 63t(20/9 +15/7) = 63t (150/t) 140t + 135t =9450<-----how do you get that? 275t= 9450 275/275 = 9450/275 t= 34 4/11 <--- how do you get that?

OpenStudy (jtvatsim):

when you say "<----- how do you get that?" do you mean just the part to the right of the equals sign, or do you mean the entire line of math?

OpenStudy (anonymous):

just that part of the equation

OpenStudy (anonymous):

meaning just this part (140t + 135t =9450)

OpenStudy (jtvatsim):

ok, that helps

OpenStudy (anonymous):

sorry about that haha

OpenStudy (jtvatsim):

So, we begin with 63t(20/9 +15/7) = 63t (150/t). On the left we distribute the 63t to get, 63t * (20/9) + 63t * (15/7) = 63t * (150/t) Does that make sense so far?

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

ok, so now just look at each piece: for 63t * (20/9) we know that 63 is just 7*9 so the 9s cancel leaving us with 7t*20 = 140t

OpenStudy (jtvatsim):

Do the same pattern for 63t * (15/7) to get 9t * 15 = 135t

OpenStudy (jtvatsim):

And then, for the right we have 63t * (150/t) but the t's cancel out! leaving us with 63*150 = 9450. Any questions on that whole deal? :)

OpenStudy (anonymous):

ohhhh okay I get it now

OpenStudy (jtvatsim):

awesome! congrats! did you still want to go over the last line of the solution?

OpenStudy (anonymous):

yes pleasee

OpenStudy (jtvatsim):

ok, so we start with the "improper fraction" 9450/275 and for some reason we want to convert this to a mixed fraction, which will be 34 4/11, but let's see how we can do that.

OpenStudy (anonymous):

okay

OpenStudy (jtvatsim):

first let's try to simplify the 9450/275 by dividing top and bottom by 5. In other words how many times does 5 go into 9450 and how many times into 275?

OpenStudy (anonymous):

1890 and 55

OpenStudy (jtvatsim):

Ok, so we've reduced the fraction to 1890/55, looks like we can still "take out" more 5s! So how many times does 5 go into 1890 and 55?

OpenStudy (jtvatsim):

by the way, we're almost there :)

OpenStudy (jtvatsim):

I get 378/11, how about you?

OpenStudy (anonymous):

I got the same

OpenStudy (jtvatsim):

alright! so we can't reduce any farther, since 11 doesn't go into 378 evenly, so now for the grand finale

OpenStudy (jtvatsim):

we will convert to a "mixed fraction" by figuring out how many 11s can fit into 378 and what remainder is left over.

OpenStudy (anonymous):

34.36

OpenStudy (jtvatsim):

OK, so we can get 11 in 34 times (ignore the .36 part)

OpenStudy (jtvatsim):

notice that 11 times 34 is 374, which means that there are 4 pieces left over to reach 378 the target number.

OpenStudy (anonymous):

What do you mean by that ?

OpenStudy (jtvatsim):

so that means that as a "mixed number" 378/11 is the same as 34 4/11, literally 34 whole pieces and 4 small pieces left over

OpenStudy (jtvatsim):

ah, you have a question? :)

OpenStudy (anonymous):

i don't get how you get the 4

OpenStudy (jtvatsim):

OK, good question.

OpenStudy (jtvatsim):

Let me use an easier example to explain.

OpenStudy (jtvatsim):

suppose we were trying to convert 15/11 into a mixed fraction.

OpenStudy (jtvatsim):

Do you see that for this easier case that the answer should be 1 4/11? If not, no worries, I just want to see exactly where the confusion might be :)

OpenStudy (anonymous):

no haha

OpenStudy (anonymous):

I am really bad in math

OpenStudy (jtvatsim):

Oh no, not at all, it's not obvious. :) Don't feel bad. No one is "good" or "bad" at math. We either see it one way or not that way. :)

OpenStudy (anonymous):

haha that just made my day

OpenStudy (jtvatsim):

lol, glad to help! :) Let's see if we can make sense of this weird symbol 15/11.

OpenStudy (jtvatsim):

First, though, would you agree that 11/11 is the same as 1?

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

OK, so then does it seem right that 15/11 = 11/11 + 4/11 ?

OpenStudy (anonymous):

yes

OpenStudy (jtvatsim):

Alright, so we just said that 11/11 is the same as 1, so I will write this as 15/11 = 1 + 4/11

OpenStudy (jtvatsim):

Now, people are lazy, and so the shorthand was invented to write 15/11 = 1 4/11 and they just made the "plus" sign disappear for no good reason! But that's the way we've written it ever since! Crazy history!

OpenStudy (jtvatsim):

And this is called "mixed fractions," which is really just adding fractions together! Any questions on this? :)

OpenStudy (anonymous):

Nope I get it now. :D

OpenStudy (jtvatsim):

Alright, so if you felt "bad" at math, it was really the inventors of math who did something "bad" by confusing people when they removed the plus sign!

OpenStudy (jtvatsim):

So, going way back to our original 378/11...

OpenStudy (jtvatsim):

We could write this as 378/11 = 11/11 + 11/11 + 11/11 + 11/11 + ... (34 times) and have 4/11 left over

OpenStudy (anonymous):

ahhh okay

OpenStudy (jtvatsim):

How do I know that? Well, I know that 378/11 = 34.36, and the whole number is 34, so that's a clue that 11 * 34 = 374

OpenStudy (jtvatsim):

In other words, 378/11 = 374/11 + 4/11

OpenStudy (jtvatsim):

but 374/11 = 34 a whole number! yay!

OpenStudy (jtvatsim):

so 378/11 = 34 + 4/11 or 34 4/11 in shorthand

OpenStudy (jtvatsim):

crazy math people! :D they like to be tricky! :)

OpenStudy (anonymous):

Yes very tricky haha && yayy ! I finally get it now haha thankt you so much^-^ you just saved me 4 hours hahah

OpenStudy (jtvatsim):

Glad to help, I've wasted many an hour on math problems myself, so I know the feeling, you did great! If you understood the explanation, you are actually very good at math! Maybe you just haven't read or heard the best explanations up to this point. Keep up the good work! :D

OpenStudy (anonymous):

Thank you :)

OpenStudy (jtvatsim):

Take care, it's lunch time here so I'm outta here!

OpenStudy (anonymous):

Okay haha have a good lunch

OpenStudy (jtvatsim):

thx